# A Time Series Analysis of Assertions in the Linux Kernel

### v1

@@ Abstract.
Assertions are a classical and typical software development technique. These are extensively used also in operating systems and their kernels, including the Linux kernel. The paper patches a gap in existing knowledge by empirically examining the longitudinal evolution of assertion use in the Linux kernel. According to the results, the use of assertions that cause a kernel panic has slightly but not substantially decreased from the kernel’s third to the sixth release series. At the same time, however, the use of softer assertion variants has increased; these do not cause a panic by default but instead produce warnings. With these time series results, the paper contributes to the existing but limited empirical knowledge base about operating system kernels and their long-term evolution.
Design by contract; kernel panic; software evolution; subsystems
ccs: Software and its engineering Operating systems
1. Introduction
@@
Assertions are a classical programming technique used to validate a program’s execution at runtime. They also serve to document code for other developers. Assertions are typically used within functions to test preconditions placed at the beginning of a function and postconditions placed at the end of the function, although constraints can be placed also upon return values and intermediate states of the function (Rosenblum95, ). In software engineering and the computing world in general, assertions are also behind a classical debate between so-called defensive programming and design by contract.
@@
In general, the defensive variant seeks to particularly validate each and every input, whereas the latter assumes an implicit contract made explicit through assertions between a caller and a supplier. In terms of the design by contract paradigm, a supplier can assume that certain conditions are always guaranteed by a caller, whereas the caller can assume a safe output provided by the supplier. When operating with assertions instead of defensive programming practices, “any runtime violation of an assertion is not a special case but always the manifestation of a software bug”, tracing of which can be started by examining the given preconditions, implying “a bug in the client (caller)”, and the given postconditions associated with “a bug in the supplier (routine)” (Meyer92, , p. 44). Regardless whichever side one takes in this debate, assertions are frequently used—and triggered—also in operating systems and their kernels. This background motives the paper’s focus on the Linux kernel.
@@
With respect to the Linux kernel, assertions have frequently been observed to be behind many of the crashes triggered by continuous kernel fuzzing frameworks (Ruohonen19RSDA, ). In the kernel context an explicitly triggered crash corresponds with a kernel panic. Without an assertion causing a panic, an alternative behavior might be that a whole operating system just hangs. Also these cases have historically been common in the Linux kernel (Gu03, ; Yoshimura12, ). Thus, the design by contract paradigm has a slightly different interpretation in the kernel context within which failing spectacularly may often, or at least sometimes, be a better option than to continue execution in undefined waters (FreeBSD24, ). Different fallback and recovery solutions are often also difficult to design and implement in the kernel context.
@@
These and related points motivated a lively debate recently among kernel developers and other open source developers. The motivating argument put forward was that Linux kernel developers should slowly move away from using assertions that cause an outright panic; instead, they should prefer warnings and try to come up with means to recover from fatal errors (LWN24a, ). In the Linux kernel the former panic-inducing assertions are used with different BUG()-style macros, whereas the latter are typically used with different WARN()-style variants. For a lack of a better terminology, these are hereafter known as “hard” and “soft” assertions. Regarding the hard assertions, the kernel documentation argues that “‘I’m too lazy to do error handling’ is not an excuse for using BUG()”, and that major “internal corruptions with no way of continuing may still use BUG(), but need good justification” (Kernel24a, ). With this guidance and the debate in mind, two research questions (RQs) are examined:
@@ 2. Related Work
In addition to the works already referenced, a couple of larger knowledge bases are worth mentioning. These allow to also better motivate the RQs. Thus, regarding RQ.1, the software evolution research domain provides a good reference point. According to the so-called laws of software evolution, which were developed by Lehman and associates from the 1970s onward (Lehman97, ), many—if not the most—software artifacts are continuously changing, continuously growing, and continuously getting more complex (Ruohonen15JSEP, ). These laws correlate with the more recent concept of technical debt; it is necessary to pay back maintenance debt in order to continue to have a given software artifact in a reasonable state in the face of continuous growth and complexity. The Linux kernel has been seen as a prime example supporting the Lehman’s laws (Israeli10, ). Regarding growth, particularly new device drivers continuously added have been the driving force (Chou01, ; Ruohonen24EASE, ). Partially due to the growth, the Linux kernel is also continuously and heavily refactored, as manifested by the extensive amount of code churn (Ruohonen19RSDA, ). In other words, technical debt is also continuously paid back. These points provide a motivation also for RQ.1. Although developers are encouraged to move away from hard assertions, retaining the existing but perhaps implicit design by contract principle requires that also soft assertions should grow alongside the growth of the kernel’s code base.
@@
The second knowledge base is the somewhat limited amount of existing empirical research on assertions. This claim should not be interpreted to mean that empirical research would be absent; particularly the benefits of assertions in software testing have been investigated empirically (Taromirad23, ). In this regard, assertions have been observed to correlate with the effectiveness of test suites (Zhang15, ) and less buggy files (Kudrjavets06, ), among other things. Another empirical research area has focused on the benefits of assertions for troubleshooting and debugging, understandability, and documentation (Patel17, ; Peruma24, ). However, to the best of the author’s knowledge, no empirical studies have been conducted about the longitudinal evolution of the use of assertion in the Linux kernel or any other kernel for that matter. Thus, the paper fills a small but notable gap in existing knowledge.
3. Research Setup
3.1. Data and Time Series
@@
The Linux kernel more or less follows the traditional semantic versioning scheme. Release series are the highest level of abstraction; currently, the Linux kernel operates with the sixth release series. Major releases, such as $6.1,6.2,6.3,\ldots$, are made based on a loosely time-based release process; a new major release occurs roughly every two or three months (Kernel24b, ). The adjective major signifies that such releases may contain everything from new features to internal changes in the kernel’s application programming interface. Each major release typically changes several hundred thousand lines of code (LOC). In addition, each major release is maintained by a separate team who manages the release engineering of minor releases, such as $6.1.1,6.1.2,6.1.3$, and so forth. These minor releases typically contain only bug fixes. Therefore, the major releases are more suitable for observing the evolution of assertions in the kernel.
@@
Thus, each major release in the third, fourth, fifth, and sixth release series was downloaded from the kernel’s online archives (Kernel24c, ). At the time of writing (20 December 2024), the amounts of major releases released in these release series were 20, 21, 20, and 13, respectively. Thus, the length of all time series constructed is 74 releases. Although the length is rather short in general, it is still sufficient for time series analysis with formal methodologies.
@@
Both the overall software architecture of the Linux kernel and its software development rely on different subsystems (Ruohonen24EASE, ). In what follows, the time series analysis is also based on major subsystems of the kernel. The major ones observed are: arch (architectural code dependent on instruction set architectures), drivers (device drivers), fs (file systems), net (networking), and sound (sound architectures). By following existing research (JiangJiang24, ; Ruohonen24EASE, ), these were proxied by the top-level directories in the kernel’s file system representation. In addition, a category of others was used to merge other notable subsystems, such as kernel for general kernel code, crypto for cryptographic implementations, and security for the kernel’s security implementations. In addition, this leftover category contains all other top-level directories insofar as they contained C programming code, as identified by files ending to a .c suffix.
@@
Table 1. Average Sizes of the Subsystems Across Releases
@@
The computations are carried out with the plain $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$ time series as well scaled variants of the two time series. Two simple scaling functions are used: the first scales the plain series by the total amount of hard and soft assertions in the $i$:th subsystems of the $t$:th release, and the second by a subsystem’s size at a given major release. In other words:
(1) $\overline{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{BUG}_{it}+\textmd{WARN}_{it}}\quad\textmd{and}\quad\widetilde{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{FILE}_{it}},$
@@
and likewise for $\textmd{WARN}_{it}$. The scaling by $\textmd{FILE}_{it}$ is arbitrary; none of the results reported changed when scaled by $\textmd{LOC}_{it}$. Finally, it can be remarked that a replication package is available online (Ruohonen24rep, ).
3.2. Methods
@@
Statistical methods for cumulative sum (CUSUM) processes are suitable for examining the research questions. The history behind CUSUM processes is rich, starting from the 1950s and peaking in the 1970s during which first formal statistical tests were derived for testing structural changes in time series (Brown75, ). By a structural change it is meant that a model estimated for a time series does not remain stable over time. The concept is much broader than the concept of (covariance) stationary, which only requires that a mean and a variance of a time series remain constant. However, in the present context these two concepts are closely related because the focus is on the means of $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$. With respect to RQ.2, a weak requirement is that the mean of the former should not remain constant. A strong requirement implies that the means of both series should not remain constant because developers are encouraged to move from hard assertions to the softer variants.
@@
The CUSUM processes are thus computed for both “static mean models” and “dynamic autoregression models”, both estimated with standard ordinary least squares (OLS). To clarify these modeling concepts, the former models merely estimate $\textmd{BUG}_{it}=\alpha_{t}+\epsilon_{t}$ and $\textmd{WARN}_{it}=\beta_{t}+\varepsilon_{t}$, where $\alpha_{t}$ and $\beta_{t}$ are intercepts, and the residuals $\epsilon_{t}$ and $\varepsilon_{t}$ are assumed to have zero means and variances $\sigma^{2}_{\epsilon_{t}}$ and $\sigma^{2}_{\varepsilon_{t}}$ for $t=1,\ldots,74$. The underlying (null) hypotheses are that $\alpha_{1}=\alpha_{2}=\cdots=\alpha_{74}$ and $\sigma^{2}_{\epsilon_{1}}=\sigma^{2}_{\epsilon_{2}}=\cdots=\sigma^{2}_{\epsilon_{74}}$, and likewise for $\beta_{t}$ and $\sigma^{2}_{\varepsilon_{t}}$. In particular, the conditional means should remain constant, although the word conditional is a slight misnomer in the present context because only constants are included in the models.
@@
The CUSUM processes are particularly useful as visual diagnostic tools. The strucchange package (Zeileis02, ) for R is used for the visualizations and associated computations. It allows to visualize the CUSUM processes computed with visual boundaries; 99% confidence level boundaries are used for the visualizations presented. In addition, the package offers more formal so-called M-fluctuation tests (Zeileis06, ). These are used to compute analogous models than the static (OLS) mean models used for the visualizations except that a Poisson regression is used for the plain $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$ time series. The four scaled series $\overline{\textmd{BUG}}_{it}$, $\widetilde{\textmd{BUG}}_{it}$, $\overline{\textmd{WARN}}_{it}$, and $\widetilde{\textmd{WARN}}_{it}$ are tested with OLS because they do not represent count data any more.
In addition, to examine RQ.1 in a little more detail, a simple non-linear growth curve estimation is carried out. Although numerous growth curves have been specified in the literature (Ruohonen15COSE, ), the standard logistic growth curve is sufficient for the present purposes. It is written the growthcurver package (Sprouffske16, ) used for computation as:
(2) $y_{t}=\frac{\alpha}{1+\frac{\alpha-y_{0}}{y_{0}}e^{-\beta t}},$
@@ 4. Results
Figure 1. $\textmd{FILE}_{it}$, $\textmd{BUG}_{it}$, and $\textmd{WARN}_{it}$
Table 2. Logistic Growth Rate Parameters and Their Standard Errors (SEs)1
@@
Table 3. Generalized M-fluctuation Tests ($p$-values)
@@
Subsystem arch $0.0003$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
drivers $0.0002$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
fs $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
net $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
sound $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
others $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $0.0001$
Figure 2. OLS-CUSUM Processes for $\textmd{FILE}_{it}$ and $\textmd{BUG}_{it}$ (the circled boundaries denote a $p<0.001$ confidence level)
Figure 3. OLS-CUSUM Processes for AR(1) models of $\Delta\textmd{FILE}_{it}$ and $\Delta\textmd{BUG}_{it}$ (the boundaries denote a $p<0.001$ confidence level)
With these observations about growing sizes of most subsystems, the logistic growth curve estimates can be subsequently considered. The growth rate parameters estimated are thus shown in Table 2. As was noted in the previous Section 3.2, some of the estimates are not reasonable, but regarding those that are, it is important to emphasize that the coefficients for $\textmd{WARN}_{it}$ are in the majority of subsystems well in line with those estimated for the two time series $\textmd{FILE}_{it}$ and $\textmd{LOC}_{t}$ measuring software size. In fact, the growth rates of soft assertion use have in some subsystems surpassed the growth rates of software size. Furthermore, in the fs subsystem and particularly in the net subsystem the ratio $\textmd{WARN}_{it}~{}/~{}\textmd{FILE}_{it}$ has exceeded a value one. In general, the answer to RQ.1 is positive; in most major subsystems the growing sizes have been met with equally growing uses of soft assertions. Having said that, it remains a subject of a debate whether the growth rates of $\textmd{WARN}_{it}$ should be even higher in order to satisfy the ever-growing sizes of some subsystems, the implicit design by contract principle, and the overall goal of providing non-fatal debugging information to users.
@@
As can be seen from Fig. 3, the static mean models for the $\textmd{BUG}_{it}$ time series mostly stay within the circled boundaries in the arch and drivers subsystems. Only minor boundary crossings are present for these two time series. In all other subsystems the boundaries are heavily crossed by the static mean models for both $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$. Although there is no right or wrong answer to a question whether one should prefer the visual inspection over the formal M-fluctuation tests, it can be concluded that the answer to RQ.2 is not entirely uniform. Given the question’s phrasing “some or all”, the conclusion is on the side of the former word. This conclusion warrants a brief speculation on what might explain the stable use of hard assertions in the two outlying subsystems across time. A reasonable presumption might be that both arch and drivers operate close to hardware, and thus implementing recovery routines may be particularly difficult in these subsystems. There may also be bugs in hardware, and existing observations indicate that many Linux device drivers tolerate such bugs poorly (Kadav09, ). The drivers subsystem has also been observed to be the most bug-prone one, including with respect to regression bugs (Chou01, ; Ruohonen24EASE, ). To that end, it might also be that hard assertions are used in the device driver subsystem to annotate code against verified and fixed bugs.

### v2

@@ Abstract
Assertions are a classical and typical software development technique. These are extensively used also in operating systems and their kernels, including the Linux kernel. The paper fills a gap in existing knowledge by empirically examining the longitudinal evolution of assertion use in the Linux kernel. According to the results, the use of assertions that cause a kernel panic has slightly but not substantially decreased from the kernel’s third to the sixth release series. At the same time, the use of softer assertion variants has increased; these do not cause a panic by default but instead produce warnings. With these time series results, the paper contributes to the existing but limited empirical knowledge base about operating system kernels and their long-term evolution.
Keywords:
D
esign by contract, kernel panic, software evolution, subsystems
1 Introduction
@@
Assertions are a classical programming technique used to verify a program’s execution at runtime. Their history traces all the way back to the early works by von Neumann and Turing [7]. In addition to verification, assertions serve many other functions too. Among other things, they also serve to document code for other developers, which can be particularly valuable when operating with legacy systems and code [11]. Assertions are typically used within functions to test preconditions placed at the beginning of a function and postconditions placed at the end of the function, although constraints can be placed also upon return values and intermediate states of the function [23]. Assertions can be used also for testing many other conditions, such as whether a condition or a state is satisfied or reached at some time in the future [1, 7, 36]. In software engineering and the computing world in general, assertions are also behind a classical debate between so-called defensive programming and design by contract.
@@
In general, the defensive variant does not trust anything coming from outside, thus seeking to particularly validate each and every input [31]. In contrast, the design by contract programming philosophy assumes an implicit contract made explicit through assertions between a caller and a supplier; a supplier can assume that certain conditions are always guaranteed by a caller, whereas the caller can assume a safe output provided by the supplier. Some implementations have made the implicit contract visible by using a keyword assume for precondition checks and a keyword promise for postcondition checks [22]. Given such checks, “any runtime violation of an assertion is not a special case but always the manifestation of a software bug”, tracing of which can be started by examining the given preconditions, implying “a bug in the client (caller)”, and the given postconditions associated with “a bug in the supplier (routine)” [17, p. 44]. Regardless whichever side one takes in this classical debate, assertions are frequently used—and triggered—also in operating systems and their kernels. The use is presumably explained partially by the context: given the overhead caused by defensive programming practices particularly for resource-constrained systems, assertions have the obvious benefit that these can be turned off during compilation [7]. This background motives the paper’s focus on the Linux kernel.
@@
With respect to the Linux kernel, assertions have frequently been observed to be behind many of crashes triggered by kernel fuzzing frameworks [28]. In the kernel context an explicitly triggered crash corresponds with a kernel panic. Without an assertion causing a panic, an alternative behavior might be that a whole operating system just hangs. Also these cases have historically been common in the Linux kernel [10, 37]. Thus, the design by contract philosophy has a slightly different interpretation in the kernel context within which failing spectacularly may often, or at least sometimes, be a better option than to continue execution in undefined waters [2, 11]. Different fallback and recovery solutions are often also difficult to design and implement in the kernel context.
@@
These and related points motivated a lively debate recently among kernel developers and other open source developers. The motivating argument put forward was that Linux kernel developers should slowly move away from using assertions that cause an outright panic; instead, they should prefer warnings and try to come up with means to recover from fatal errors [8]. In the Linux kernel the former panic-inducing assertions are used with different BUG()-style macros, whereas the latter are typically used with different WARN()-style variants. For a lack of a better terminology, these are hereafter known as “hard” and “soft” assertions. Regarding the hard assertions, the kernel documentation argues that “‘I’m too lazy to do error handling’ is not an excuse for using BUG()”, and that major “internal corruptions with no way of continuing may still use BUG(), but need good justification” [33]. With this guidance and the debate in mind, the following two research questions (RQs) are examined:
@@ 2 Related Work
In addition to the works already referenced, a couple of larger knowledge bases are worth mentioning. These allow to also better motivate the RQs. Thus, regarding RQ.1, the software evolution research domain provides a good reference point. According to the so-called laws of software evolution, which were developed by Lehman and associates from the 1970s onward [16], many—if not the most—software artifacts are continuously changing, continuously growing, and continuously getting more complex [27]. These laws correlate with the more recent concept of technical debt; it is necessary to pay back maintenance debt in order to continue to have a given software artifact in a reasonable state in the face of continuous growth and increasing complexity. The Linux kernel has been seen as a prime example supporting the Lehman’s laws [12]. Regarding growth, particularly new device drivers continuously added have been the driving force [6, 25]. Partially due to the growth, the Linux kernel is also continuously and heavily refactored, as manifested by the extensive amount of code churn [28]. In other words, technical debt is also continuously paid back. These points provide a motivation also for RQ.1. Although developers are encouraged to move away from hard assertions, retaining the existing but perhaps implicit design by contract principle requires that also soft assertions should grow alongside the growth of the kernel’s code base.
@@
The second knowledge base is the somewhat limited amount of existing empirical research on assertions. This claim should not be interpreted to mean that empirical research would be absent altogether; among other things, the benefits of assertions in software testing have been investigated empirically [5, 30]. In this regard, assertions have been observed to correlate with the effectiveness of test suites [40] and less buggy files [15]. Another empirical research area has focused on the benefits of assertions for troubleshooting and debugging, reusability, understandability, and documentation [18, 20, 21, 35]. This research area aligns closely with the design by contract philosophy. However, to the best of the author’s knowledge, no empirical studies have been conducted about the longitudinal evolution of the use of assertion in the Linux kernel—or any other operating system kernel for that matter. Thus, the paper fills a small but notable gap in existing software engineering knowledge.
3 Research Setup
3.1 Data and Time Series
@@
The Linux kernel more or less follows the traditional semantic versioning scheme. Release series are the highest level of abstraction; currently, the Linux kernel operates with the sixth release series. Major releases, such as $6.1,6.2,6.3,\ldots$, are made based on a loosely time-based release process; a new major release occurs roughly every two or three months [32]. The adjective major signifies that such releases may contain everything from new features to internal changes in the kernel’s application programming interface. Each major release typically changes several hundred thousand lines of code (LOC). In addition, each major release is maintained by a separate team who manages the release engineering of minor releases, such as $6.1.1,6.1.2,6.1.3$, and so forth. These minor releases typically contain only bug fixes. Therefore, the major releases are more suitable for observing the evolution of assertions in the kernel.
@@
Thus, each major release in the third, fourth, fifth, and sixth release series was downloaded from the kernel’s online archives [34]. At the time of writing (20 December 2024), the amounts of major releases released in these release series were 20, 21, 20, and 13, respectively. Thus, the length of all time series constructed is 74 releases. Although the length is rather short in general, it is still sufficient for time series analysis with formal methodologies.
@@
Both the overall software architecture of the Linux kernel and its software development rely on different subsystems [25]. In what follows, the time series analysis is also based on major subsystems of the kernel. The major ones observed are: arch (architectural code dependent on instruction set architectures), drivers (device drivers), fs (file systems), net (networking), and sound (sound architectures). By following existing research [13, 25], these were proxied by the top-level directories in the kernel’s file system representation. In addition, a category of others was used to merge other notable subsystems, such as kernel for general kernel code, crypto for cryptographic implementations, and security for the kernel’s security implementations. In addition, this leftover category contains all other top-level directories insofar as they contained C programming code, as identified by files ending to a .c suffix.
@@
Table 1: Average Sizes of the Subsystems Across Releases
@@
To give a short example about the use of the assertion variants in practice, the file ./drivers/base/regmap/regcache.c triggers a kernel panic in a function in case a precondition is not satisfied by a caller:
        void regcache_exit(struct regmap *map)
                if (map->cache_type == REGCACHE_NONE)
                        return;
                BUG_ON(!map->cache_ops);
The above case is a good example because it could be argued that a softer assertion variant might work or that an explicit check could be done analogously to checking for the REGCACHE_NONE constant. The same file uses also BUG() as a default fallback case in a switch statement, indicating for a caller that only the values explicitly listed in the statement are assumed to be valid.
@@
Then, the computations are carried out with the plain $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$ time series as well scaled variants of the two time series. Two simple scaling functions are used: the first scales the plain series by the total amount of hard and soft assertions in the $i$:th subsystems of the $t$:th release, and the second by a subsystem’s size at a given major release. In other words:
$\overline{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{BUG}_{it}+\textmd{WARN}_{it}}\quad\textmd{and}\quad\widetilde{\textmd{BUG}}_{it}=\frac{\textmd{BUG}_{it}}{\textmd{FILE}_{it}},$ (1)
@@
and likewise for $\textmd{WARN}_{it}$. The scaling by $\textmd{FILE}_{it}$ is arbitrary; none of the results reported changed when scaled by $\textmd{LOC}_{it}$. Finally, it can be remarked that a replication package is available online [24]. It contains the dataset and the code.
3.2 Methods
@@
Statistical methods for cumulative sum (CUSUM) processes are suitable for examining the research questions. The history behind CUSUM processes is rich, starting from the 1950s and peaking in the 1970s during which first formal statistical tests were derived for testing structural changes in time series [3]. By a structural change it is meant that a model estimated for a time series does not remain stable over time. To this end, a term parameter stability is sometimes used as an alternative [4]. The concept is much broader than the concept of (covariance) stationary, which only requires that a mean and a variance of a time series remain constant. However, in the present context these two concepts are closely related because the focus is on the means of $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$. With respect to RQ.2, a weak requirement for a positive answer is that the means of the former should not remain constant for all or at least most $i$. A strong requirement implies that the means of both series should not remain constant in all or most subsystems because developers are encouraged to move from hard assertions to the softer variants.
@@
The CUSUM processes are thus computed for both “static mean models” and “dynamic autoregression models”, both estimated with standard ordinary least squares (OLS). To clarify these modeling concepts, the former models merely estimate $\textmd{BUG}_{it}=\alpha_{t}+\epsilon_{t}$ and $\textmd{WARN}_{it}=\beta_{t}+\varepsilon_{t}$, where $\alpha_{t}$ and $\beta_{t}$ are intercepts, and the residuals $\epsilon_{t}$ and $\varepsilon_{t}$ are assumed to have zero means and variances $\sigma^{2}_{\epsilon_{t}}$ and $\sigma^{2}_{\varepsilon_{t}}$ for $t=1,\ldots,74$. The underlying (null) hypotheses are that $\alpha_{1}=\alpha_{2}=\cdots=\alpha_{74}$ and $\sigma^{2}_{\epsilon_{1}}=\sigma^{2}_{\epsilon_{2}}=\cdots=\sigma^{2}_{\epsilon_{74}}$, and likewise for $\beta_{t}$ and $\sigma^{2}_{\varepsilon_{t}}$. In particular, the conditional means should remain constant, although the word conditional is a slight misnomer in the present context because only constants are included.
@@
The CUSUM processes are particularly useful as visual diagnostic tools. The strucchange package [39] for R is used for the visualizations and associated computations. It allows to visualize the CUSUM processes computed with visual boundaries; 99% confidence level boundaries are used for the visualizations presented. With the earlier points in mind and using the eventually assertion notation proposed in the literature [1], the weak requirement for a positive answer to RQ.2 means that eventually("the boundaries are crossed" $\equiv$ TRUE) for all or at least most $i$. The logic is thus similar to using CUSUM processes for anomaly detection [19]. In addition, the strucchange package offers more formal so-called M-fluctuation tests [38]. These are used to compute analogous models than the static (OLS) mean models used for the visualizations except that a Poisson regression is used for the plain $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$ time series. The four scaled series $\overline{\textmd{BUG}}_{it}$, $\widetilde{\textmd{BUG}}_{it}$, $\overline{\textmd{WARN}}_{it}$, and $\widetilde{\textmd{WARN}}_{it}$ are tested with OLS because they do not represent count data any more.
In addition, to examine RQ.1 in a little more detail, a simple non-linear growth curve estimation is carried out. Although numerous growth curves have been specified in the literature [26], the standard logistic growth curve is sufficient for the present purposes. It is written in the growthcurver package [29] used for computation as:
$y_{t}=\frac{\alpha}{1+\frac{\alpha-y_{0}}{y_{0}}e^{-\beta t}},$ (2)
@@ 4 Results
Figure 1: $\textmd{FILE}_{it}$, $\textmd{BUG}_{it}$, and $\textmd{WARN}_{it}$
Table 2: Logistic Growth Rate Parameters and Their Standard Errors (SEs)1
@@
Table 3: Generalized M-fluctuation Tests ($p$-values)
@@
Subsystem arch $0.0003$ <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
drivers $0.0002$ <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
fs <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
net <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
sound <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001
others <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 <0.0001<0.0001 $0.0001$
Figure 2: OLS-CUSUM Processes for $\textmd{FILE}_{it}$ and $\textmd{BUG}_{it}$ (the circled boundaries denote a $p<0.001$ confidence level)
Figure 3: OLS-CUSUM Processes for AR(1) models of $\Delta\textmd{FILE}_{it}$ and $\Delta\textmd{BUG}_{it}$ (the boundaries denote a $p<0.001$ confidence level)
With these observations about growing sizes of most subsystems, the logistic growth curve estimates can be subsequently considered. The growth rate parameters estimated are thus shown in Table 2. As was noted in the previous Section 3.2, some of the estimates are not reasonable, but regarding those that are, it is important to emphasize that the coefficients for $\textmd{WARN}_{it}$ are in the majority of subsystems well in line with those estimated for the two time series $\textmd{FILE}_{it}$ and $\textmd{LOC}_{t}$ measuring software size. In fact, the growth rates of soft assertion use have in some subsystems surpassed the growth rates of software size. Furthermore, in the fs subsystem and particularly in the net subsystem the ratio $\textmd{WARN}_{it}~/~\textmd{FILE}_{it}$ has exceeded a value one. In general, the answer to RQ.1 is positive; in most major subsystems the growing sizes have been met with equally growing uses of soft assertions. Having said that, it remains a subject of a debate whether the growth rates of $\textmd{WARN}_{it}$ should be even higher in order to satisfy the ever-growing sizes of some subsystems, the implicit design by contract principle, and the overall goal of providing non-fatal debugging information.
@@
As can be seen from Fig. 3, the static mean models for the $\textmd{BUG}_{it}$ time series mostly stay within the circled boundaries in the arch and drivers subsystems. Only minor boundary crossings are present for these two time series. In all other subsystems the boundaries are heavily crossed by the static mean models for both $\textmd{BUG}_{it}$ and $\textmd{WARN}_{it}$. Although there is no right or wrong answer to a question whether one should prefer the visual inspection over the formal M-fluctuation tests, it can be concluded that the answer to RQ.2 is not entirely uniform. Given the question’s phrasing “some or all”, the conclusion is on the side of the former word. This conclusion warrants a brief speculation on what might explain the stable use of hard assertions in the two outlying subsystems across time. A reasonable presumption might be that both arch and drivers operate close to hardware, and thus implementing recovery routines may be particularly difficult in these subsystems. There may also be bugs in hardware, and existing observations indicate that many Linux device drivers tolerate such bugs poorly [14]. The drivers subsystem has also been observed to be the most bug-prone one, including with respect to regression bugs [6, 25]. To that end, it might also be that hard assertions are used in the device driver subsystem to annotate code against verified and fixed bugs.